OpenStudy (anonymous):

Convert 0.0000001 to a power of 10. 107 10−7 106 10−6

OpenStudy (anonymous):

Any help would be sweggy

OpenStudy (anonymous):

@acxbox22 @helptheblond

OpenStudy (acxbox22):

;)

OpenStudy (acxbox22):

since it is less than 0 we know the power of ten has to be less than 1

OpenStudy (acxbox22):

because 10^1=10 and 10^0=1

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

The answers are 10 to a power of 7 10to a power of −7 10to a power of 6 10to a power of −6

OpenStudy (acxbox22):

from the info above we can cancel 10^7 and 10^6

OpenStudy (anonymous):

move the dot until you get 10 and since you are moving to the right, you receive a power of -1 for each time you move a digit to the right

jimthompson5910 (jim_thompson5910):

Place the decimal point after the first nonzero digit (which is 1) to get 1.0 How many places do you have to move that decimal point to get back to 0.0000001 ?

OpenStudy (anonymous):

7 ?

jimthompson5910 (jim_thompson5910):

correct, you have to move the decimal point 7 spots to the left to go from 1.0 to 0.0000001

OpenStudy (anonymous):

6 i mean

jimthompson5910 (jim_thompson5910):

so that means \[\Large 0.0000001 = 1.0 \times 10^{-7}\]

OpenStudy (acxbox22):

^^^^

jimthompson5910 (jim_thompson5910):

notice how the -7 in the exponent represents moving that decimal point 7 spots to the left

jimthompson5910 (jim_thompson5910):

and because 1 times any number = same number, we can say \[\Large 0.0000001 = 1.0 \times 10^{-7} = 10^{-7}\]

OpenStudy (anonymous):

So the answer is 10^−7 Correct ?

jimthompson5910 (jim_thompson5910):

you mean 10^(-7) or 10^-7 right?

OpenStudy (anonymous):

Yes

jimthompson5910 (jim_thompson5910):

it's correct

OpenStudy (anonymous):

Thanks

jimthompson5910 (jim_thompson5910):

\[\Large 10\wedge(-7) = 10^{-7}\]

OpenStudy (anonymous):

 do the same for all and it makes this problems simple